OpenStudy (anonymous):

Integral of x^2 e^(x^3) Integral of x^2 e^(x^3) @Mathematics

OpenStudy (lalaly):

\[\large{u=x^3}\]\[\large{du=3x^2dx}\] \[\large{\frac{1}{3} \int\limits{e^udu}} = \frac{1}{3}e^u\]substitute x^3 back instead of u \[\huge{\int\limits{x^2e^{x^3}dx}} = \frac{1}{3}e^{x^3} +c\]

OpenStudy (anonymous):

Thanks for you

OpenStudy (lalaly):

:)